A garden measuring 12 meters by 16 meters is to have a pedestrian pathway installed all around it, increasing the total area to 285 square meters. What will be the width of the pathway?
Answer by checkley77(12844)

(16+2x)(12+2x)=285
192+24x+32x+4x^2=285
4x^2+56x+192-285=0
4x^2+56x-93=0
(2x-3)(2x+31)=0
2x=3
x=3/1
x=1.5 ans. for the width of the pathway.
Proof:
(16+2*1.5)(12+2*1.5)=285
(16+3)(12+3)=285
19*15=285
285=285